OpenStudy (anonymous):

A rectangle has perimeter of 14cm and length of x. Show that the width of the rectangle is (7-x)??

Parth (parthkohli):

\[2(x + w) = 14\]\[\implies 2x + 2w = 14\]\[\implies 2w = 14 -2x\]\[\implies w = 7 - x\]

Parth (parthkohli):

Do you get the above?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Good job!

Parth (parthkohli):

For confirmation, can you tell what I did there? :)

OpenStudy (anonymous):

yeah just solved for w.

OpenStudy (anonymous):

thank you kindly

Parth (parthkohli):

You're welcome :)